OpenStudy (anonymous):

how to set this up? how many quarts of pure antifreeze must be added to 12 quarts of a 10% antifreeze solution to obtain a 40% antifreeze solution?

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

0.1(12) + x = 0.4(12 + x)

OpenStudy (anonymous):

x is amount of quarts. since it is pure it is 100% it was 0.1(12) + 1*x = 0.4(12 + x)

OpenStudy (anonymous):

Ok thanks i can do the rest from here